OpenStudy (anonymous):

what equation passes through the line of (-2,-1) and (-4,-3)

OpenStudy (anonymous):

Which of the following is the equation of a line that passes through (-2, -1) and (-4, -3)? A. y = x + 1 B. y = x + 1 C. y = x – 1 D. y = x – 1

OpenStudy (anonymous):

lol...ok

OpenStudy (hesan):

\[(x+2) \div (-4+2) = (y+1) \div (-3+1)\]

OpenStudy (hesan):

Solve it to get the eq